hello, i am having an extreme difficult time with this radical stuff. Can you help me solve the following problem:
radical X+4 + radical X+ 20= 8

Answer by stanbon(29478)

radical (X+4) + radical(X) + 20= 8
sqrt(x+4) + sqrtx = -12
Square both sides:
x+4+2sqrt(x(x+4))+x = 144
Simplify:
2sqrt(x^2+4x) = -2x+40
Divide thru by 2 to get:
sqrt(x^2+4x) = -x+20
Square both sides to get:
x^2+4x = x^2-40x+400
Simplify to get:
44x=400
x=100/11
Cheers,
Stan H.